# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ 4.6.1. Reductions
After removing at most finitely many points from $C$, we can assume that $C$ is contained in $S^{3}_{K}$ and we denote by $\mathcal{E}_{i}\to C$ the standard elliptic curve schemes on $C$. After possibly replacing $K$ with a finite extension and $\overline{C}$ with a finite étale cover, we obtain a covering datum
$\mathcal{D}:=(\mathfrak{G}^{\prime}_{1},\mathfrak{G}^{\prime}_{2},\mathfrak{G}^{\prime}_{3},\tilde{\mathfrak{C}}^{\prime\prime},\tilde{\mathfrak{C}}^{\prime},\tilde{C}^{\prime},\tilde{C},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
for $(\overline{C},s_{0})$, and we fix a Weil height $h$ on $\tilde{C}^{\prime}$. As before, it suffices for us to show that there exist constants $c_{9}$, $c_{10}$} and $\delta$ such that, for any modular point $s:=(x_{1},x_{2},x_{3})\in C$ of supersingular exponent $\delta$, and any $\tilde{s}\in\nu^{-1}(s)\in\tilde{C}$, we have
@@ 4.6.2. Setup
We copy the notations and repeat the constructions of Section 4.5, this time with $I:=\{1,2,3\}$. In particular, for any $s=(x_{1},x_{2},x_{3})$ and $\tilde{s}$, as above, and any $w\in\Sigma_{K(\tilde{s}),f}$ lying above any $v\in\Sigma_{K,f}$, the condition $|x(\tilde{s})|_{w}<R^{\leq}_{v}$ implies
$\mathfrak{G}^{\prime}_{[j,i],\mathfrak{s}(j),k_{w}}\cong\mathfrak{G}^{\prime}_{[j,i],\mathfrak{s}_{1}(j),k_{w}}$
for all $j\in\{1,\ldots,\ell\}$ and $i\in I$. It follows that, for any $i\in I$, the $j$-invariants $x_{i}$ and $b_{i}$ are congruent modulo the maximal ideal of $\mathcal{O}_{K(\tilde{s}),w}$, which is to say, $|x_{i}-b_{i}|_{w}<1$.
@@ 4.6.3. Constructing relations
$Q^{\rm nord}:=\prod_{v\in\Sigma^{\rm nord}_{\tilde{s}}}Q_{v}$
with $Q_{v}$ a relation of degree at most $2$ and $\Sigma^{\rm nord}_{\tilde{s}}\subset\Sigma_{\tilde{s}}$ the complement of the set of finite places $v\in\Sigma_{\tilde{s}}$ for which the $\mathcal{E}_{i,s}$ have good ordinary reduction. The fact that $Q$ is nontrivial is proved exactly as in Section 4.5.4, hence, we omit it.
@@
$|\Sigma^{\rm sup}_{\tilde{s}}|\leq\Delta(s)^{\delta}$
and so the result follows from [And89, Thm. 5.2] provided $3(6\ell-1)\delta<1/2$. Rearranging, this is equivalent to
$\delta<6^{-1}(6\ell-1)^{-1}$
and, since $\ell$ depends only on $C$ and $s_{0}$, the proof is complete. ∎
5. Proving Theorems 3 and 4
The proofs of Theorems 3 and 4 follow the Pila–Zannier strategy. In fact, following word-for-word the proof of [HP12, Prop. 5.1], it suffices to obtain the following large Galois orbit results, which serve as substitutes for [HP12, Lemma 4.2].
Theorem 17.
Let $C$ satisfy the conditions of Theorem 3. There exist constants $c_{12}$ and $c_{13}$ such that, for any point $(x_{1},x_{2},x_{3})$ as in the statement of Conjecture 2,
@@ 5. Proving Theorems 3 and 4
Theorem 18.
Let $C$ and $(b_{1},b_{2},b_{3})$ be as in Theorem 4. There exist constants $c_{14}$ and $c_{15}$ such that, for any point $(x_{1},x_{2},x_{3})$ as in the statement of Theorem 4,
@@
The proofs of these theorems follow almost exactly the proof of [DO22, Prop. 5.15], which is essentially an extract from the proof of [HP12, Lemma 4.2]. The key input is a Masser–Wüstholz-type isogeny estimate—we cite the (effective) bound of Gaudron–Rémond [GR14, Thm. 1.4].
Appendix: rigid analytic neighbourhoods
We establish a version of [DO22, Lemma 5.5], written in terms of rigid geometry instead of only $\mathbb{C}_{p}$-points, and thereby give a rigid analytic proof of [DO22, Lemma 5.5]. We also establish a version which talks about the neighbourhood of a single point where $x$ has a simple zero, instead of requiring all zeroes of $x$ to be simple.
First we check that an étale morphism of rigid spaces restricts to an isomorphism on some open neighbourhood of each $K$-point. (Note that it is not true in general that étale morphisms are local isomorphisms – the open neighbourhoods on which they are isomorphisms need not form an admissible covering.)
Lemma 19.
@@
Let $Z$ denote the connected component of $Y$ which contains $x_{0}$. Since $Z$ is connected and it has a $K$-point $x_{0}$, $Z_{L}$ is connected [Con99, Thm. 3.2.1]. Hence $Z_{L}$ is a connected component of $Y_{L}$, that is, $Z_{L}=Y_{i}$ for some $i$. Thus $f_{L}|_{Z_{L}}\colon Z_{L}\to\bar{D}(0,r^{\prime},L)$ is an isomorphism, so $f|_{Z}\colon Z\to\bar{D}(0,r^{\prime},k)$ is an isomorphism [Con06, Thm. A.2.4]. ∎
Proposition 20.
Let $C$ be a smooth irreducible algebraic curve over a number field $K$. Let $x\in K(C)$ be a rational function. Let $s_{0}\in C(K)$ be a simple zero of $x$.
For each $v\in\Sigma_{K,f}$, there exists a real number $r_{v}>0$ and a rigid open subspace $U_{v}\subset C^{v\textrm{-}\mathrm{an}}$ with the following properties:
@@
By standard facts about algebraic curves, $C$ is isomorphic to a Zariski open subset of some smooth projective curve over $K$. The conclusions of the proposition are local around a given point of $C$, so we may replace $C$ by this smooth projective curve.
Consider $x$ as a morphism $C\to\mathbb{P}^{1}_{K}$. By [DO23, Lemma 6.1], there exists a regular $\mathcal{O}_{K}$-model $\mathfrak{C}$ of $C$ such that $x$ extends to a morphism of $\mathcal{O}_{K}$-schemes $\xi\colon\mathfrak{C}\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$. Since $x$ is a non-constant morphism between integral curves, it is flat. Since $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ is also flat, the fibre-wise criterion of flatness establishes that $\xi\colon\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ is flat at all points of the generic fibre.
By hypothesis, $x$ is unramified, hence étale, at $s_{0}$. Therefore, by [Gro67, Rmk. 17.8.3], $\xi$ is étale at $s_{0}$. It follows that $\xi$ is étale on a non-empty Zariski open neighbourhood of $s_{0}$ in $\mathfrak{C}$.
@@
By [DO23, Lemma 6.3], there exists a section $\mathfrak{s}_{0}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathfrak{C}$ which extends $s_{0}\colon\operatorname{Spec}(K)\to C$. It follows that there exists a dense open subset $\mathfrak{V}\subset\operatorname{Spec}(\mathcal{O}_{K})$ such that $\xi$ is étale on $\mathfrak{s}_{0}(\mathfrak{V})$.
@@
We shall give two constructions of open subspaces $U_{v}\subset C^{v\textrm{-}\mathrm{an}}$ satisfying (ii) and (iii), depending on whether $\mathfrak{p}_{v}$ lies in $|\mathfrak{V}|$ or not. The construction for primes in $|\mathfrak{V}|$ gives $r_{v}=1$, while the construction for other primes may give $r_{v}<1$. This is sufficient to ensure that (i) holds.
@@
First consider $v\in\Sigma_{K,f}$ such that $\mathfrak{p}:=\mathfrak{p}_{v}\in|\mathfrak{V}|$. Let $\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}$ denote the formal completion of $\mathfrak{C}$ along $\mathfrak{s}_{0}(\mathfrak{p})$ and let $\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}}$ denote the formal completion of $\mathbb{P}^{1}_{\mathcal{O}_{K}}$ along $\mathfrak{z}_{0}(\mathfrak{p})$, where $\mathfrak{z}_{0}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$ denotes the zero section. Since $\xi$ is étale at $\mathfrak{s}_{0}(\mathfrak{p})$, it induces an isomorphism of formal schemes $\xi_{\mathfrak{p},\mathrm{for}{\dagger}}\colon\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}\to\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}}$. Applying Berthelot’s rigid generic fibre functor, we obtain an isomorphism of rigid spaces $\xi_{\mathfrak{p}}^{\mathrm{rig}}\colon\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}^{\mathrm{rig}}\to(\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}})^{\mathrm{rig}}$. Here, $\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}^{\mathrm{rig}}$ is a rigid open subspace of $C^{v\textrm{-}\mathrm{an}}$, and $(\mathbb{P}^{1}_{\mathfrak{p},\mathrm{for}{\dagger}})^{\mathrm{rig}}=\operatorname{Spf}(\mathcal{O}_{K,\mathfrak{p}}[\![X]\!],(\mathfrak{p},X))^{\mathrm{rig}}$ is the rigid open unit disc. Thus, if we define $U_{v}$ to be $\mathfrak{C}_{\mathfrak{p},\mathrm{for}{\dagger}}^{\mathrm{rig}}$, then $U_{v}$ satisfies (ii) and (iii) with $r_{v}=1$.
@@
Now consider $v\in\Sigma_{K,f}$ such that $\mathfrak{p}_{v}\notin|\mathfrak{V}|$ (this construction applies for all $v\in\Sigma_{K,f}$, but we use the previous construction for $v$ such that $\mathfrak{p}_{v}\in|\mathfrak{V}|$). Since $x$ is finite and étale on a Zariski open neighbourhood of $s_{0}$ in $C$, there is a rigid open neighbourhood $V_{v}$ of $x$ in $C^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}|_{V_{v}}$ is finite étale. After shrinking $V_{v}$, we may assume that the image of $V_{v}$ is contained in a disc around 0. By Lemma 19, there exists an open subspace $U_{v}\subset V_{v}$ containing $s_{0}^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}|_{U_{v}}$ is an isomorphism of rigid spaces from $U_{v}$ to a closed rigid disc $\bar{D}(0,r_{v},K_{v})$. After shrinking $U_{v}$, we may instead arrange that $x^{v\textrm{-}\mathrm{an}}|_{U_{v}}$ is an isomorphism onto the open rigid disc $D(0,r_{v},K_{v})$. Thus $U_{v}$ satisfies (ii) and (iii).
@@
Now we prove a rigid analogue of [DO22, Lemma 5.5].
Proposition 21.
Let $C$ be a smooth algebraic curve over a number field $K$. Let $x\in K(C)$ be a rational function of degree $\ell$, such that $x$ has $\ell$ distinct unramified zeros $s_{1},\dotsc,s_{\ell}\in C(K)$.
For each $v\in\Sigma_{K,f}$, there exists a real number $r_{v}>0$ and rigid open subspaces $U_{v,1},\dotsc,U_{v,\ell}\subset C^{v\textrm{-}\mathrm{an}}$ with the following properties:
@@
Apply Proposition 20 to each of $s_{1},\dotsc,s_{k}$, thereby yielding $r_{v,1},\dotsc,r_{v,\ell}$ and $U_{v,1},\dotsc,U_{v,\ell}$ for all $v\in\Sigma_{K,f}$. These satisfy (ii), (iv) and (v). However, for a given $v$, the subspaces $U_{v,1},\dotsc,U_{v,\ell}$ might not be pairwise disjoint.
@@
As in the proof of Proposition 20, we may choose a regular model $\mathfrak{C}$ of $C$ such that $x$ extends to a morphism $\xi\colon\mathfrak{C}\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$, as well as sections $\mathfrak{s}_{1},\dotsc,\mathfrak{s}_{k}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathfrak{C}$ which extend $s_{1},\dotsc,s_{\ell}$. There is a dense open subset $\mathfrak{V}\subset\operatorname{Spec}(\mathcal{O}_{K})$ such that $\xi$ is étale on $\mathfrak{s}_{k}(\mathfrak{V})$ for all $k=1,\dotsc,\ell$. Since $s_{1},\dotsc,s_{k}$ are pairwise distinct, after shrinking $\mathfrak{V}$, we may also assume that $\mathfrak{s}_{j}(\mathfrak{p})\neq\mathfrak{s}_{k}(\mathfrak{p})$ for all $\mathfrak{p}\in|\mathfrak{V}|$ and all $j\neq k$.
We will show that, whenever $\mathfrak{p}_{v}\in|\mathfrak{V}|$, the $U_{v,1},\dotsc,U_{v,\ell}$ will be pairwise disjoint and form an admissible covering of their union. Indeed, let $\mathfrak{C}_{\mathfrak{p}}$ denote the fibre above $\mathfrak{p}:=\mathfrak{p}_{v}$ of $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$. Let $\mathfrak{C}_{\mathfrak{p},\mathrm{for}}$ and $\mathfrak{C}_{\mathfrak{p},\mathrm{for},k{\dagger}}$ denote the formal completions of $\mathfrak{C}$ along $\mathfrak{C}_{\mathfrak{p}}$ and $\mathfrak{s}_{k}(\mathfrak{p})$ respectively. Since $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ is proper, $\mathfrak{C}_{\mathfrak{p}}^{\mathrm{rig}}\to C^{v\textrm{-}\mathrm{an}}$ is an isomorphism, so we obtain a reduction map $\operatorname{red}\colon C^{v\textrm{-}\mathrm{an}}\to\mathfrak{C}_{\mathfrak{p}}$. By construction, we have
$U_{v,k}=\mathfrak{C}_{\mathfrak{p},\mathrm{for},k{\dagger}}^{\mathrm{rig}}=\operatorname{red}^{-1}(\mathfrak{s}_{k}(\mathfrak{p})).$
@@
Since $\mathfrak{s}_{1}(\mathfrak{p}),\dotsc,\mathfrak{s}_{\ell}(\mathfrak{p})$ are distinct (by our choice of $\mathfrak{p}\in|\mathfrak{V}|$), we deduce that $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint.
Consider the sets
$Z_{\mathfrak{p},k}:=\mathfrak{C}_{\mathfrak{p}}\setminus\{s_{j}:1\leq j\leq\ell,\,j\neq k\}$

### v2

@@ 4.6.1. Reductions
After removing at most finitely many points from $C_{0}$, we can assume that $C_{0}$ is contained in $S^{3}_{K}$. Thus, after possibly replacing $K$ with a finite extension and $C$ with a finite étale cover, we obtain a covering datum
$\mathcal{D}:=(\mathfrak{G}_{1},\mathfrak{G}_{2},\mathfrak{G}_{3},\mathfrak{C}^{\prime},{\mathfrak{C}},\tilde{C},\tilde{C}_{0},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
for $(C,s_{0})$, and we fix a Weil height $h$ on $\tilde{C}$. As before, it suffices for us to show that there exist constants $c_{10}$, $c_{11}$} and $c_{12}$} such that, for any modular point $s:=(x_{1},x_{2},x_{3})\in C$ of supersingular exponent $\delta$, and any $\tilde{s}\in\nu^{-1}(s)\subset\tilde{C}_{0}$, we have
@@ 4.6.2. Setup
We copy the notations and repeat the constructions of Section 4.5.2, this time with $I:=\{1,2,3\}$. We construct a graph $G$ as in Section 4.5.3, but this time with edges between $[k,j]$ and $[k,j+1]$ for $j\in\{1,2\}$. Thus, to each $e\in E(G)$, we assign this $j:=j(e)$ and, then, e(i):=[k(e),j(e)+(i1)]e(i):=[k(e),j(e)+(i-1)]. Copying the rest of Section 4.5.3, we obtain an isogeny datum
$\{C,\mathfrak{C},\mathfrak{G}_{e(i)},C_{0},\mathcal{E}_{e(i),0},\mathfrak{C}_{0},s_{1},f_{e},\{\omega_{e(i)},\eta_{e(i)}\},x,F_{e(i)mn},R_{v},\Delta_{v}:1\leq i,m,n\leq 2,v\in\Sigma_{K}\}$
for each $e\in E(G)$.
@@ 4.6.3. Constructing relations
Let $s\in C_{0}$ with $\nu(s)=(x_{1},x_{2},x_{3})$ modular of supersingular exponent $\delta$. Replace $K$ with $K(s)$ and write $\mathfrak{s}_{k}$ and $\mathfrak{s}$ for the $\mathcal{O}_{K}$-extensions of $s_{k}$ and $s$, respectively, in $\mathfrak{C}^{\prime}$. Note that, for any $v\in\Sigma_{K}$, the condition $|x(s)|_{v}<R_{v}$ implies $s^{v\textrm{-}\mathrm{an}}\in U_{R_{v},k}$ for some $k\in\{1,\ldots,\ell\}$. If $v\in\Sigma_{K,f}$, then, since $U_{R_{v},k}$ is a simple neighbourbood, we have
$\mathfrak{G}_{i,\mathfrak{s},k_{v}}\cong\mathfrak{G}_{i,\mathfrak{s}_{k},k_{v}}$
for all $i\in I$, which implies that the $j$-invariants $x_{i}$ and $b_{i}$ are congruent modulo the maximal ideal of $\mathcal{O}_{K_{v}}$, which is to say, $|x_{i}-b_{i}|_{v}<1$.
Define $\Sigma_{s}\subset\Sigma_{K}$ to be the set of $v\in\Sigma_{K}$ satisfying $|x(s)|_{v}<R_{v}$. In particular, $\Sigma_{s}\cap\Sigma_{K,f}$ is contained in the set of $v\in\Sigma_{K,f}$ for which (2) holds. By repeating the arguments of Section 4.5, invoking Proposition 3.6 in the place of Proposition 3.2, we obtain a global relation between the evaluations at $x(s)$ of the elements of $\mathcal{G}_{H}$, which is equal to the product of a relation of degree at most $4$ (dealing with the good ordinary places) and a relation
$Q^{\rm nord}:=\prod_{v\in\Sigma^{\rm nord}_{s}}Q_{v}$
where each $Q_{v}$ is a relation of degree at most $2$ and $\Sigma^{\rm nord}_{s}\subset\Sigma_{s}$ is the complement in $\Sigma_{s}$ of the set of finite places for which the $\mathcal{E}_{i,s}$ have good ordinary reduction. The fact that $Q$ is nontrivial is proved exactly as in Section 4.5.6, hence, we omit it.
4.6.4. Constraining the degree
@@
$|\Sigma^{\rm sup}_{s}|\leq[K:\mathbb{Q}]\Delta(s)^{\delta}$
and so the result follows from [And89, Thm. 5.2]. ∎
Remark 4.3.
From [And89, VII, Thm. 5.2], noting that the $\overline{\mathbb{Q}}(X)$-linear span of $\mathcal{G}_{H}$ is closed under differentiation, we see that Theorem 1.8 holds for any
@@
5. Proving Theorems 1.3 and 1.6
The proofs of Theorems 1.3 and 1.6 follow the Pila–Zannier strategy. In fact, following word-for-word the proof of [HP12, Prop. 5.1], it suffices to obtain the following large Galois orbit results, which serve as substitutes for [HP12, Lemma 4.2].
Theorem 5.1.
Let $C$ satisfy the conditions of Theorem 1.3. There exist constants $c_{14},c_{15}>0$ such that, for any modular point $s=(x_{1},x_{2},x_{3})\in C$, we have
@@
Theorem 5.2.
[…]
Appendix A Rigid analytic neighbourhoods
We establish a version of [DO25, Lemma 5.5], written in terms of rigid geometry instead of only $\mathbb{C}_{p}$-points, and thereby give a rigid analytic proof of [DO25, Lemma 5.5]. Along the way, we establish a version which talks about the neighbourhood of a single point where $x$ has a simple zero, instead of requiring all zeroes of $x$ to be simple. For convenience, we also prove that the neighbourhoods we construct can be chosen to be simple, as defined in Section 2.8.
First we check that an étale morphism of rigid spaces restricts to an isomorphism on some open neighbourhood of each $K$-point. (Note that it is not true in general that étale morphisms are local isomorphisms—the open neighbourhoods on which they are isomorphisms need not form an admissible covering.)
Lemma A.1.
@@ Appendix A Rigid analytic neighbourhoods
Let $Z$ denote the connected component of $Y$ which contains $x_{0}$. Since $Z$ is connected and it has a $K$-point $x_{0}$, $Z_{L}$ is connected [Con99, Thm. 3.2.1]. Hence $Z_{L}$ is a connected component of $Y_{L}$, that is, $Z_{L}=Y_{i}$ for some $i$. Thus $f_{L}|_{Z_{L}}\colon Z_{L}\to\bar{D}(0,r^{\prime},L)$ is an isomorphism, so $f|_{Z}\colon Z\to\bar{D}(0,r^{\prime},k)$ is an isomorphism [Con06a, Thm. A.2.4]. ∎
Henceforth, let $C$ denote a smooth projective geometrically irreducible algebraic curve over a number field $K$, and let $x\in K(C)$ be a rational function. Let $\mathfrak{C}$ be a regular projective $\mathcal{O}_{K}$-model of $C$.
Proposition A.2.
Suppose that $s_{0}\in C(K)$ is a simple zero of $x$.
For each $v\in\Sigma_{K,f}$, there exists a real number $r_{v}>0$ and a simple open subspace $U_{v}\subset C^{v\textrm{-}\mathrm{an}}$ with the following properties:
@@
Since $x\colon C\to\mathbb{P}^{1}_{K}$ is a non-constant morphism between integral curves, it is flat. By hypothesis, $x$ is unramified at $s_{0}$. Thus, $x$ is étale at $s_{0}$.
@@
By [DO23, Lem. 6.2], there exists a regular projective $\mathcal{O}_{K}$-model $\mathfrak{C}^{\prime}$ and a morphism $f:\mathfrak{C}^{\prime}\to\mathfrak{C}$ such that $x$ extends to a morphism $\xi:\mathfrak{C}^{\prime}\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$. Since $\mathfrak{C}^{\prime}\to\operatorname{Spec}(\mathcal{O}_{K})$ is flat (by the definition of an $\mathcal{O}_{K}$-model), $\xi$ is étale at $s_{0}$ by [Gro67, Rmk. 17.8.3]. It follows that $\xi$ is étale on a non-empty Zariski open neighbourhood of $s_{0}$ in $\mathfrak{C}^{\prime}$. Hence, there exists a dense open subset $\mathfrak{V}\subset\operatorname{Spec}(\mathcal{O}_{K})$ such that $\xi$ is étale on $\mathfrak{s}_{0}(\mathfrak{V})$, where $\mathfrak{s}_{0}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathfrak{C}^{\prime}$ is the $\mathcal{O}_{K}$-extension of $s_{0}$.
@@
We shall give two constructions of simple open subspaces $U_{v}\subset C^{v\textrm{-}\mathrm{an}}$ satisfying (ii) and (iii), depending on whether $\mathfrak{p}_{v}$ lies in $|\mathfrak{V}|$ or not. The construction for primes in $|\mathfrak{V}|$ gives $r_{v}=1$, while the construction for other primes may give $r_{v}<1$. This is sufficient to ensure that (i) holds.
First consider $v\in\Sigma_{K,f}$ such that $\mathfrak{p}:=\mathfrak{p}_{v}\in|\mathfrak{V}|$. Let $\mathfrak{C}^{\prime}_{\mathfrak{p}}$ denote the fibre above $\mathfrak{p}:=\mathfrak{p}_{v}$ of $\mathfrak{C}^{\prime}\to\operatorname{Spec}(\mathcal{O}_{K})$. Let $\hat{\mathfrak{C}^{\prime}}_{\mathfrak{C}^{\prime}_{\mathfrak{p}}}$ and $\hat{\mathfrak{C}^{\prime}}_{\mathfrak{s}_{0}(\mathfrak{p})}$ denote the formal completions of $\mathfrak{C}^{\prime}$ along $\mathfrak{C}^{\prime}_{\mathfrak{p}}$ and $\mathfrak{s}_{0}(\mathfrak{p})$ respectively. Define $\hat{\mathfrak{C}}_{\mathfrak{C}_{\mathfrak{p}}}$ analogously. Since $\mathfrak{C}^{\prime}\to\operatorname{Spec}(\mathcal{O}_{K})$ and $\mathfrak{C}\to\operatorname{Spec}(\mathcal{O}_{K})$ are proper, the open immersions $\hat{\mathfrak{C}^{\prime}}_{\mathfrak{C}^{\prime}_{\mathfrak{p}}}^{\mathrm{rig}}\to\hat{\mathfrak{C}}_{\mathfrak{C}_{\mathfrak{p}}}\to C^{v\textrm{-}\mathrm{an}}$ are isomorphisms.
Let $U_{v}:=(\hat{\mathfrak{C}}^{\prime})_{\mathfrak{s}_{0}(\mathfrak{p})}^{\mathrm{rig}}$. By [Ber96, Prop. 0.2.7], $U_{v}$ is a simple open subspace of $C^{v\textrm{-}\mathrm{an}}$ with respect to $\mathfrak{C}^{\prime}$. The existence of the following commutative diagram implies that $U_{v}$ is a simple open subspace of $C^{v\textrm{-}\mathrm{an}}$ with respect to $\mathfrak{C}$:
[…]
Indeed, it is immediate that $U_{v}$ satisfies properties (1) and (2) in the definition of simple open subspace with respect to $\mathfrak{C}$. Property (3) follows from the fact that the value of $\operatorname{red}_{v}$ on $U$ is equal to the image of $f(\mathfrak{s}_{0}(\mathfrak{p}))$ in $\mathfrak{C}_{k_{v}}$, which is a smooth point by [Liu06, Exercise 4.3.25(c)].
@@
Let $\hat{\mathbb{P}}^{1}_{\mathfrak{z}_{0}(\mathfrak{p})}$ denote the formal completion of $\mathbb{P}^{1}_{\mathcal{O}_{K}}$ along $\mathfrak{z}_{0}(\mathfrak{p})$, where $\mathfrak{z}_{0}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$ denotes the zero section. Since $\xi$ is étale at $\mathfrak{s}_{0}(\mathfrak{p})$, it induces an isomorphism of formal schemes $\hat{\xi}_{\mathfrak{s}_{0}(\mathfrak{p})}\colon\hat{\mathfrak{C}}^{\prime}_{\mathfrak{s}_{0}(\mathfrak{p})}\to\hat{\mathbb{P}}^{1}_{\mathfrak{z}_{0}(\mathfrak{p})}$. Applying Berthelot’s rigid generic fibre functor, we obtain an isomorphism of rigid spaces $\hat{\xi}_{\mathfrak{s}_{0}(\mathfrak{p})}^{\mathrm{rig}}\colon U_{v}=(\hat{\mathfrak{C}}^{\prime})_{\mathfrak{s}_{0}(\mathfrak{p})}^{\mathrm{rig}}\to\hat{\mathbb{P}}^{1,\mathrm{rig}}_{\mathfrak{z}_{0}(\mathfrak{p})}$. Since $\hat{\mathbb{P}}^{1,\mathrm{rig}}_{\mathfrak{z}_{0}(\mathfrak{p})}=\operatorname{Spf}(\mathcal{O}_{K,\mathfrak{p}}[\![X]\!],(\mathfrak{p},X))^{\mathrm{rig}}$ is the rigid open unit disc, we conclude that $U_{v}$ satisfies (iii) with $r_{v}=1$. Clearly, $U_{v}$ satisfies (ii).
@@
Now consider $v\in\Sigma_{K,f}$ such that $\mathfrak{p}_{v}\notin|\mathfrak{V}|$ (this construction applies for all $v\in\Sigma_{K,f}$, but we use the previous construction for $v$ such that $\mathfrak{p}_{v}\in|\mathfrak{V}|$). Since $x$ is finite and étale on a Zariski open neighbourhood of $s_{0}$ in $C$, there is a rigid open neighbourhood $V_{v}$ of $s^{v\textrm{-}\mathrm{an}}_{0}$ in $C^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}|_{V_{v}}$ is finite étale. After shrinking $V_{v}$, we may assume that the image of $V_{v}$ is contained in a disc around 0. By Lemma A.1, there exists an open subspace $U_{v}\subset V_{v}$ containing $s_{0}^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}|_{U_{v}}$ is an isomorphism of rigid spaces from $U_{v}$ to a closed rigid disc $\bar{D}(0,r_{v},K_{v})$. After shrinking $U_{v}$ and $r_{v}$, we may arrange that $U_{v}$ is simple and that $x^{v\textrm{-}\mathrm{an}}|_{U_{v}}$ is instead an isomorphism onto the open rigid disc $D(0,r_{v},K_{v})$. Thus, $U_{v}$ satisfies (ii) and (iii).
@@
Now we prove a rigid analogue of [DO25, Lemma 5.5].
Proposition A.3.
Suppose that $x$ has $\ell$ distinct unramified zeros $s_{1},\dotsc,s_{\ell}\in C(K)$.
For each $v\in\Sigma_{K,f}$, there exists a real number $r_{v}>0$ and simple open subspaces $U_{v,1},\dotsc,U_{v,\ell}\subset C^{v\textrm{-}\mathrm{an}}$ with the following properties:
@@
Apply Proposition A.2 to each of $s_{1},\dotsc,s_{\ell}$, thereby yielding $r_{v,1},\dotsc,r_{v,\ell}$ and $U_{v,1},\dotsc,U_{v,\ell}$ for all $v\in\Sigma_{K,f}$. The latter automatically satisfy (ii), (iv) and (v). However, they need not satisfy (iii).
@@
By [DO23, Lem. 6.2], there exists a regular projective $\mathcal{O}_{K}$-model $\mathfrak{C}^{\prime}$ and a morphism $f:\mathfrak{C}^{\prime}\to\mathfrak{C}$ such that $x$ extends to a morphism $\xi:\mathfrak{C}^{\prime}\to\mathbb{P}^{1}_{\mathcal{O}_{K}}$. Let $\mathfrak{s}_{1},\dotsc,\mathfrak{s}_{k}\colon\operatorname{Spec}(\mathcal{O}_{K})\to\mathfrak{C}^{\prime}$ denote the $\mathcal{O}_{K}$-extensions of $s_{1},\ldots,s_{\ell}$. As in the proof of Proposition A.2, there is a dense open subset $\mathfrak{V}\subset\operatorname{Spec}(\mathcal{O}_{K})$ such that $\xi$ is étale on $\mathfrak{s}_{k}(\mathfrak{V})$ for all $k=1,\dotsc,\ell$. Since $s_{1},\dotsc,s_{k}$ are pairwise distinct, after shrinking $\mathfrak{V}$, we may also assume that $\mathfrak{s}_{j}(\mathfrak{p})\neq\mathfrak{s}_{k}(\mathfrak{p})$ for all $\mathfrak{p}\in|\mathfrak{V}|$ and all $j\neq k$.
Now, suppose first that $\mathfrak{p}_{v}\in|\mathfrak{V}|$. As in the proof of Proposition A.2, we can and do suppose that
$U_{v,k}=(\hat{\mathfrak{C}^{\prime}})_{\mathfrak{s}_{k}(\mathfrak{p})}^{\mathrm{rig}}=(\operatorname{red}_{v}^{\prime})^{-1}(\mathfrak{s}_{k}(\mathfrak{p}))$
@@
and $r_{v}:=r_{v,k}=1$ for all $k=1,\ldots,\ell$. Since $\mathfrak{s}_{1}(\mathfrak{p}),\dotsc,\mathfrak{s}_{\ell}(\mathfrak{p})$ are distinct (by our choice of $\mathfrak{p}\in|\mathfrak{V}|$), we deduce that $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint.
Now consider the sets
$Z_{\mathfrak{p},k}:=\mathfrak{C}^{\prime}_{\mathfrak{p}}\setminus\{\mathfrak{s}_{j}(\mathfrak{p}):1\leq j\leq\ell,\,j\neq k\}$